### v1

@@ 5.1 Image-Based Baselines
We introduce a dataset comprising multi-view images and evaluate its fidelity using 2D feature embedding-based anomaly detection methodologies, specifically FastFlow [71], STPM [72], CFA [73], cflow [29] and CDO [74]. 2D reconstruction-based anomaly detection methodologies, specifically FAVAE [75] and UniAD [21]. These methods leverage neural networks to learn representations conducive to accurately reconstructing normal samples, with anomalies detected based on deviations in the reconstruction quality. Our findings are summarized in Table. 3. Additionally, we investigate the effectiveness of PAD [8] and Splat [10], a 3D multi-view reconstruction-based algorithm. Notably, both 2D and 3D reconstruction-based methods exhibit limitations. The discrepancy is primarily attributed to misalignment between the training and testing sets, compounded by variations in lighting conditions that impact image quality. Our visualization is also consistent with our findings shown in Figure. 9.

### v2

@@ 5.1 Image-Based Baselines
We introduce a dataset comprising multi-view images and evaluate its fidelity using 2D feature embedding-based anomaly detection methodologies, specifically FastFlow [74], STPM [75], CFA [76], cflow [32] and CDO [78]. 2D reconstruction-based anomaly detection methodologies, specifically FAVAE [77] and UniAD [24]. These methods leverage neural networks to learn representations conducive to accurately reconstructing normal samples, with anomalies detected based on deviations in the reconstruction quality. Our findings are summarized in Table. 3 and Table.  4. Additionally, we investigate the effectiveness of PAD [11] and Splat [14], a 3D multi-view reconstruction-based algorithm. Notably, both 2D and 3D reconstruction-based methods exhibit limitations. The discrepancy is primarily attributed to misalignment between the training and testing sets, compounded by variations in lighting conditions that impact image quality. Our visualization is also consistent with our findings shown in Figure. 9.